# Solving Distributed Flexible Job Shop Scheduling Problems in the Wool Textile Industry with Quantum Annealing

### v1

@@ Abstract
Many modern manufacturing companies have evolved from a single production site to a multi-factory production environment that must handle both geographically dispersed production orders and their multi-site production steps. The availability of a range of machines in different locations capable of performing the same operation and shipping times between factories have transformed planning systems from the classic Job Shop Scheduling Problem (JSSP) to Distributed Flexible Job Shop Scheduling Problem (DFJSP). As a result, the complexity of production planning has increased significantly. In our work, we use Quantum Annealing (QA) to solve the DFJSP. In addition to the assignment of production orders to production sites, the assignment of production steps to production sites also takes place. This requirement is based on a real use case of a wool textile manufacturer. To investigate the applicability of this method to large problem instances, problems ranging from 50 variables up to 250 variables, the largest problem that could be embedded into a D-Wave quantum annealer Quantum Processing Unit (QPU), are formulated and solved. Special attention is dedicated to the determination of the Lagrange parameters of the Quadratic Unconstrained Binary Optimization (QUBO) model and the QPU configuration parameters, as these factors can significantly impact solution quality. The obtained solutions are compared to solutions obtained by Simulated Annealing (SA), both in terms of solution quality and calculation time. The results demonstrate that QA has the potential to solve large problem instances specific to the industry.
@@ 1 Introduction
This paper is based on a use case of a textile manufacturing company that operates in the worsted yarn industry and specializes in pure wool and wool blends for weaving, flat and circular knitting, luxury, as well as the non-apparel market. This business-to-business company has production sites in several countries on different continents. Production sites in different locations perform only specific production steps and work together in sequence-dependent setups to fulfill each individual sales order. For this purpose, sales orders first become production orders that must be assigned to the various production sites for the processing of various production steps. Given the scattering of production sites that might be involved in the processing of one single production order, a thorough production planning is critical.
The reason is that this production planning problem is an NP-hard combinatorial optimization problem, which can be reduced to the Distributed Flexible Job Shop Scheduling Problem (DFJSP) in this use case. It is widely known as one of the most challenging combinatorial optimization problems [15]. At the same time, operational production planning requires continuous and process-parallel optimization of machine allocation and processing sequence.
Quantum annealing (QA) is a promising technology to solve such difficult combinatorial optimization problems. While gate-based quantum computers can handle a broader range of problems, quantum annealers are specifically designed to solve combinatorial optimization problems commonly encountered in industry [26]. In this work, we use the AdvantageSystem4.1 quantum annealer from D-Wave to solve the DFJSP. To the best of our knowledge, there is no other work that specifically addresses solving the DFJSP using QA. Our contributions are as follows:
@@
We use a QA to solve the DFJSP. Due to the specifics of the wool textile industry, we also extend the DFJSP and make it even more complex: The classic DFJSP is extended to not only consider the distribution of production orders, but also the distribution of production steps (of a production order). Shipping times between production steps at different sites must also be included in the model. To the best of our knowledge, there is no other work in literature that discusses the solution of the classical or the extended DFJSP with QA.
@@
Calculation of Lagrange parameters and QPU configuration parameters. The values of the so-called Lagrange parameters of the QUBO model and the QPU configuration parameters have a significant influence on the solution quality. Instead of the trial-and-error method, as some approaches suggest111https://www.dwavesys.com/media/qvbjrzgg/guide-2.pdf, we determine these parameters based on the mathematical formulation of the problem.
@@
Assessment from an economic perspective. We address the question whether the use of QA can lead to a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
This paper is organized as follows: In Section 2 we review the state of the art on optimization problems and quantum annealing before we discuss related work in Section 3. We describe our QUBO model to solve the extended DFJSP in Section 4. Based on the QUBO model, we describe and discuss our experiments in Section 5. Section 6 provides a summary.
@@ 2 Preliminaries
In Subsection 2.1, we first study computationally complex combinatorial optimization problems that are typical in production planning. We then turn to the topic of quantum annealing in Subsection 2.2 and to simulated annealing in Subsection 2.3. The connection between Binary Quadratic Model (BQM) / Quadratic Unconstrained Binary Optimization (QUBO) and QA is highlighted in Subsection 2.4.
@@ 2.1 Combinatorial Optimization Problems
The Job Shop Scheduling Problem (JSSP) is a combinatorial optimization problem222Combinatorial optimization is the search process for maxima (or minima) of an objective function where the domain has a discrete but large configuration space. where a given set of jobs (or production orders) must be executed while only a limited number of machines can perform the operations of the jobs (also called production steps). Each operation can be executed by a specific machine. A schedule must be found that determines which jobs are to be executed at which time and by which machine. The aim is to minimize the total amount of time until all jobs have been completed [4, 1]. Many variations of the JSSP problem exist. To be closer, both to the market and manufacturing resources, modern enterprises have evolved from a single factory to a multi-factory manufacturing environment. This has led to the scheduling system changing from a JSSP system to a Flexible Job Shop Scheduling Problem (FJSP) and further to a Distributed Flexible Job Shop Scheduling Problem (DFJSP).
The FJSP is a generalization of the basic JSSP that allows to process tasks on one machine, chosen from a set of alternative machines (the machines in each set are identical) [19]. The FJSP is considered one of the most difficult combinatorial optimization problems, and it was revealed that it is strongly NP-hard since the classical JSSP has also proven to be strongly NP-hard [28]. A distinctive feature of JSSP (and FJSP) is the parallel job processing. While different process chains of an order are valuated, the choice of a process chain for an order affects the choice of production routes for all other orders.
The DFJSP refers to production activities that are carried out by several geographically distributed factories. The DFJSP can be regarded as an extension of FJSP which also includes the selection of suitable factories for the jobs [15]. The DFJSP described in the literature implies that if a job is assigned to a factory, then all the operations of this job must be processed at the same factory [13, 15]. The wool textile industry is a special case of DFJSP, where operation of a job can be processed in different locations. Assigning specific operations to different factories results in unique production schedules for the jobs, which influences the production planning. The DFJSP is more complex than the standard FJSP and is also classified as NP-Hard in complexity theory [15]. The goal of scheduling to assign each operation to factories and machines, considers the shipping time between locations, and launches operations on machines in a way that minimizes the makespan of all operations, thus ensuring that all jobs are completed in the shortest time possible.
@@ 2.2 Quantum Annealing
Today, two main architectures for quantum computers exist: gate model (also known as circuit model) and QA [4]. Gate model quantum computers use quantum gates to perform calculations, analogous to Boolean gates in classical computers. These processors are not yet fault-tolerant or large enough to reach quantum supremacy, a state that is referred to as the Noisy Intermediate-Scale Quantum (NISQ) era [3]. Consequently, it is difficult to develop practical applications on universal quantum gate model computers at present.
On the other hand, QA devices, like the computer developed by D-Wave, map a problem to a mathematical model of interacting magnets, then search for an optimal solution for problems with a large solution space. In general, QA is a mechanical metaheuristic to solve problems with a large, but discrete, search space.
QA systems are initialized in a low-energy state. So, the parameters of a problem are changed slowly, which allows the system to end in a low-energy state of the problem that corresponds to an optimal solution [5]. The evolution of the system can be represented by a time-dependent Hamiltonian333A classical Hamiltonian is a mathematical description of some physical system in terms of its energies [5].:
where A and B are time-dependent coefficients, s is the annealing time rescaled to take values between 0 and 1, are the Pauli Matrices444In quantum mechanics, Pauli Matrices are a set of 2 × 2 complex Hermitian and unitary matrices that arise in Pauli’s treatment of spins. of operations on a qubit q, h is the qubit bias and J the coupling strength between qubits. The algorithm gradually transitions in time from the initial ground state of the Initial Hamiltonian to a state described by the Final Hamiltonian [21]. The Final Hamiltonian matches the energy function to be minimized, and represents the best solution at its minimum value.
@@ 2.3 Simulated Annealing
Simulated Annealing (SA) is a nature inspired stochastic metaheuristic, single solution-based algorithm used to find a global optimum in a large problem space [22], and can be seen as a classical computing pendant to QA. Fig. 1 shows a comparison between QA and SA.
In SA, to simulate the annealing process, a temperature variable is set high and then the system is allowed to cool while the algorithm runs. SA is using thermal effects to converge to a low energy state, but therefore needs to jump over higher energy barriers. On the other hand, QA uses the quantum tunneling effect555In quantum mechanics, the tunneling effect describes the penetration of particles through barriers, even if particle energy is less than the barrier height [8]., which allows it to potentially surpass this higher barrier directly and thus reach the minimum faster. Recent advancements in QA methods have suggested potential benefits by employing QA over SA to address NP-hard optimization challenges [16].
@@ 2.4 Binary Quadratic Models
For QA, cost functions are mathematical expressions of the problem formulated as quadratic models that have the lowest values for good solutions. To solve problems on D-Wave quantum annealer, the problem space has to be formulated as a BQM, the D-Wave Systems software representation of the objective function:
where s and x are binary variables, i.e. categorical variables that can only take one of two values. s is the decision variable called spin that can either take values from the set {-1,1}, in which case the problem is called an Ising problem (Eq. 2) [14]. In the Ising model, h is the magnetic field that represents the linear coefficients corresponding to qubit biases and J represents the coupling strength between qubits [27]. In a QUBO problem (Eq. 3), x can take values from the set {0,1} [12]. Qubit biases are represented by the linear coefficients a and the coupling strength between qubits by the quadratic coefficients b. Both versions are NP-hard and can easily be converted into each other by applying variable substitution [25]. In this paper, the QUBO format is used to formulate the problem.
To solve a problem on a D-wave quantum annealer, the problem is formulated as an objective function in QUBO format. Low energy states of the objective function represent good solutions of the problem. Because the cost function can be represented as a graph, it can be mapped onto the QPU: linear coefficients to qubit biases and quadratic coefficients to coupler strengths. The QPU uses quantum annealing to seek the minimum in the resulting energy landscape, which corresponds to the solution of the problem. QA devices are less affected by noise than gate model quantum computers. This allows more qubit usage, and thus solves larger problems. QA provides an approach that promises good solutions to NP-hard problems and is specialized in combinatorial optimization. While the discussion continues whether QA can provide quantum speed-up over classical approaches, there have been several attempts to solve application-motivated problems with QA [26].
@@ 3 Related Work
In this Section, we assess previous approaches to solve the optimization problems described in Subsection 2.1.
@@
JSSP. In the past a variety of approaches to solve the JSSP or its variation existed. Denkena et al. [9] mention heuristics, Genetic Algorithm (GA), Tabu Search, Ant Colony, Particle Swarm Optimization and SA. These approaches are used both individually and in hybrid combinations with one another. Venturelli et al. [24] present a solver for the JSSP that makes use of a QA. Carugno et al. [4] study the application of QA to solve JSSP and discuss challenges related to problem formulation to the fine-tuning of the QA. Carugno et al. [4] study the application of QA to solve the JSSP. In terms of solution quality, QA remarkably effective in comparison to classical approaches on very constrained problem instances, while closely competitive on realistic problem instances. Pakhomchik et al. [20] focus on the mathematical formulation of JSSP and propose a new large neighborhood search approach, with good empirical results for large problem instances. Zielewski et al. [29] use QA to solve JSSP and especially investigate the embedding process in the performance. The authors show that the initial setup of the problem, post-processing and problem formulation, as well as excess variables and large embeddings have a significant impact on results. A. Sharma and P. Maharjan [2] compare QA to SA through existing published works and conclude that QA has the potential to be considered as an optimized technique against SA. QA is more efficient than SA in estimating the ground-state of the system for small and for large sample size simulations. Denchev et al. [8] show that, for problems with tall and narrow energy barriers separating local minima, QA is more than times faster than SA running on a single core.
@@
Overall, research considering transportation time focusses on flow shop scheduling problem666Flow shop scheduling is a special case of job shop scheduling in which all jobs have identical operation flow patterns on machines [1]. or classical FJSP. Most of the QA research in this field focuses on solving JSSP or FJSP and transportation time is often neglected in literature. However, transportation time between locations can have a big impact on maximum completion time of FJSP.
@@
FSP. Flow Shop Scheduling Problem (FSP) is one of the most studied problems in order planning. Hurink and Knust [10] apply SA to minimize both total completion time and total tardiness of flow shop problems. Naderi et al. [17] propose an adaptation of the SA algorithm to solve the flow shop problem with incorporated sequence-dependent transportation times with a single-transporter. Several heuristic algorithms were developed considering the transportation of an interrupted job from a machine to another machine. Naderi et al. [18] consider the transportation times in permutation flow shops and provide solution methods including heuristics and metaheuristics to solve large-sized problems. An adaptation of the imperialist competitive algorithm hybridized by a simulated annealing based local search was proposed by Karimi et al. [11] to solve the problem.
@@
FJSP. Schworm et al. [21] investigate a QA-based FJSP on large problem instances specific to modern manufacturing and demonstrate the efficiency of the approach regarding scalability, solutions’ quality, and computing time. Denkena et al. [9] compare the computing time of several approaches to solve the classical FJSP and conclude that it generally increases significantly with the problem size. Their research suggests that these methods deliver valid results on small problem instances, but are not suitable for problem sizes relevant to practice. Therefore, methods are required that yield high quality outcomes, while they also ensure consistently quick computation times, particularly for large problem instances. The paper additionally investigates the concept of QA based optimization for a process parallel FJSP on existing benchmarks. The results reveal that the QA technique consistently achieves low anneal times, especially for large instances, while it still produces outcomes comparable to those of classical heuristics.
DFJSP. Only few papers focus on the FJSP [28] with transportation time between jobs, also known as DFJSP. Zhang et al. [28] develop a modified Genetic Algorithm (GA) to solve the FJSP with transportation time. Marzouki et al. [15] use a Chemical Reaction Optimization Metaheuristic to solve the DFJSP. Liu et al. [13] propose a refined GA and apply it with satisfactory results on a real case of a manufacturing company.
@@
Generally, most of the papers focus on studying the JSSP. Although there is some research that includes transportation time in the model, these use classical approaches to solve the problem, like heuristics or GA. In this research, an extended case of DFJSP is studied, where not only jobs, but also job operations can be distributed. For this purpose, a D-Wave Quantum Annealer is used to solve the DFJSP with distributed operations. The availability of a set of machines capable to process the same operation and the geographical distribution of these machines add complexity and pose a significant challenge to solve this combinatorial optimization problem.

### v3

@@ Abstract
Many modern manufacturing companies have evolved from a single production facility to a multi-factory production environment that must manage both regionally dispersed production orders and their multi-site production steps. The availability of a range of machines in different locations capable of performing the same operation and shipping times between factories have transformed planning systems from the classic Job Shop Scheduling Problem (JSSP) to the Distributed Flexible Job Shop Scheduling Problem (DFJSP). Consequently, the complexity of production planning has increased significantly. We employ Quantum Annealing (QA) to solve the DFJSP in our research. In addition to assigning production orders to production sites, production steps are also assigned to these sites. This requirement is based on a real use case of a wool textile manufacturing company. To investigate the applicability of this method to large problem instances, problems ranging from 50 variables up to 250 variables, the largest problem that could be embedded into a D-Wave quantum annealer Quantum Processing Unit (QPU), are formulated and solved. Special attention is dedicated to determining the Lagrange parameters of the Quadratic Unconstrained Binary Optimization (QUBO) model and the QPU configuration parameters, as these factors can significantly impact solution quality. The obtained solutions are compared to solutions obtained by Simulated Annealing (SA), both in terms of solution quality and calculation time. The results demonstrate that QA has the potential to solve large problem instances specific to the industry.
@@ 1 Introduction
This paper is based on a use case of a textile manufacturing company that operates in the worsted yarn industry, specializing in pure wool and wool blends for weaving, flat and circular knitting, as well as luxury and non-apparel markets. This business-to-business company has production sites in several countries on different continents. Production sites in different locations perform only specific production steps and work together in sequence-dependent setups to fulfill each individual sales order. For this purpose, sales orders are first converted into production orders, which must be assigned to the various production sites for processing through different production steps. Given the scattering of production sites that may be involved in processing a single production order, thorough production planning is crucial.
The reason is that this production planning problem is an NP-hard combinatorial optimization problem, which can be reduced to the Distributed Flexible Job Shop Scheduling Problem (DFJSP) in this use case. It is widely known as one of the most difficult combinatorial optimization problems [17]. At the same time, operational production planning requires continuous and process-parallel optimization of machine allocation and processing sequence.
Quantum annealing (QA) is a promising technology to solve combinatorial optimization problems. While gate-based quantum computers can handle a broader range of problems, quantum annealers are specifically designed to solve combinatorial optimization problems commonly encountered in industry [36]. In this work, we use the Advantage System 4.1 quantum annealer from D-Wave to solve the DFJSP.
To the best of our knowledge, there is no other work that explicitly addresses solving the DFJSP using QA. Our three contributions are as follows:
@@
We use a QA to solve the DFJSP. Due to the specifics of the wool textile industry, we also extend the DFJSP and make it even more complex: The classical DFJSP is extended to not only consider the distribution of production orders, but also the distribution of production steps (of a production order). Shipping times between production steps at different sites must also be included in the model. To the best of our knowledge, there is no other work in the literature that discusses the solution of the classical or the extended DFJSP with QA.
@@
Calculation of Lagrange parameters and QPU configuration parameters. The values of the so-called Lagrange parameters of the QUBO model and the QPU configuration parameters have a significant influence on the solution quality. Instead of the trial-and-error method, as some approaches suggest111Programming the D-Wave QPU [White Paper]: https://www.dwavesys.com/media/qvbjrzgg/guide-2.pdf, we determine these parameters based on the mathematical formulation of the problem.
@@
Assessment from an economic perspective. We address the question of whether the use of QA can lead to a speed advantage for the DFJSP with distributed operations specific to the wool textile industry.
This paper is organized as follows: In Section 2, we review the state of the art on optimization problems and quantum annealing, before we discuss related work in Section 3. We describe our QUBO model to solve the extended DFJSP in Section 4. Based on the QUBO model, we describe and discuss our experiments in Section 5. Section 6 provides a summary.
@@ 2 Preliminaries
In Subsection 2.1, we initially study computationally complex combinatorial optimization problems that are typical in production planning. We then turn to the topic of quantum annealing in Subsection 2.2 and to simulated annealing in Subsection 2.3. The connection between the Binary Quadratic Model (BQM) / Quadratic Unconstrained Binary Optimization (QUBO) and QA is highlighted in Subsection 2.4.
@@ 2.1 Combinatorial Optimization Problems
The Job Shop Scheduling Problem (JSSP) is a combinatorial optimization problem222Combinatorial optimization is the search process for maxima (or minima) of an objective function where the domain has a discrete but large configuration space. where a given set of jobs (or production orders) must be executed. At the same time, only a limited number of machines can perform the operations required for the jobs (also referred to as production steps). Each operation can be executed by a specific machine. A schedule must be found that determines which jobs are to be executed at which time and by which machine. The aim is to minimize the total amount of time until all jobs have been completed [5, 1]. Many variations of the JSSP problem exist. To be closer to both the market and manufacturing resources, modern enterprises have evolved from a single factory to a multi-factory manufacturing environment. This has led to the scheduling system changing from a JSSP system to a Flexible Job Shop Scheduling Problem (FJSP) and further to a Distributed Flexible Job Shop Scheduling Problem (DFJSP).
The FJSP is a generalization of the JSSP that allows tasks to be processed on one machine, chosen from a set of alternative machines (the machines in each set are identical) [21]. As is well known, this is a strongly NP-hard problem [39]. A distinctive feature of JSSP (and FJSP) is the parallel job processing. While different process chains for an order are evaluated, the selection of a process chain for an order influences the choice of production routes for all other orders.
The DFJSP refers to production activities that are carried out by several geographically distributed factories. The DFJSP can be regarded as an extension of FJSP, which also includes the selection of appropriate factories for the jobs [14]. The DFJSP described in the literature implies that if a job is assigned to a factory, then all the operations of this job must be processed at the same factory [14, 17]. The wool textile industry is a special case of DFJSP, where a job’s operation can be processed in different locations. Assigning specific operations to different factories results in unique production schedules for the jobs, which influences the production planning. The DFJSP is more complex than FJSP and is also classified as NP-Hard in complexity theory [17]. The goal of scheduling is to assign each operation to factories and machines, considering the shipping time between locations, and launching operations on machines in a way that minimizes the makespan of all operations, thus ensuring that all jobs are completed in the shortest time possible.
@@ 2.2 Quantum Annealing
Today, two main architectures for quantum computers exist: gate-based quantum computers and QA [5]. Gate-based quantum computers use quantum gates to perform calculations, analogous to Boolean gates in classical computers. These processors are not yet fault-tolerant or large enough to achieve quantum supremacy, a state referred to as the Noisy Intermediate-Scale Quantum (NISQ) era [4]. Consequently, it is difficult to develop practical applications on gate-based quantum computers at present.
On the other hand, QA devices, such as the computer developed by D-Wave, map a problem to a mathematical model of interacting magnets and then search for an optimal solution for problems with a large solution space. Si et al. [27] developed a large-scale and energy-efficient superparamagnetic Ising quantum annealer, which advances the scope of coherent Ising devices for combinatorial optimization. In general, QA is a mechanical metaheuristic to solve problems with a large, but discrete, search space.
QA systems are initialized in a low-energy state. The problem parameters are changed slowly enough, which allows the system to end in a low-energy state of the problem that corresponds to an optimal solution [36]. The evolution of the system can be represented by a time-dependent Hamiltonian333See also Carugno et al. [5] and https://docs.dwavequantum.com/en/latest/quantum_research/quantum_annealing_intro.html for the Hamiltonian term.:
where A and B are time-dependent coefficients, s is the annealing time rescaled to take values between 0 and 1, are the Pauli Matrices444In quantum mechanics, Pauli Matrices are a set of 2 × 2 complex Hermitian and unitary matrices that arise in Pauli’s treatment of spins. of operations on a qubit qi, hi is the qubit bias, and Ji,j the coupling strength between qubits. The algorithm gradually transitions in time from the initial ground state of the Initial Hamiltonian to a state described by the Final Hamiltonian [25]. The Final Hamiltonian matches the energy function to be minimized, and represents the best solution at its minimum value.
@@ 2.3 Simulated Annealing
Simulated Annealing (SA) is a nature-inspired stochastic metaheuristic, single solution-based algorithm used to find a global optimum in a large problem space [29], and can be seen as a classical computing pendant to QA. Fig. 1 shows a comparison between QA and SA.
In SA, to simulate the annealing process, a temperature variable is set high, and then the system is allowed to cool while the algorithm runs. SA is using thermal effects to converge to a low energy state, but therefore needs to jump over higher energy barriers. On the other hand, QA uses the quantum tunneling effect555In quantum mechanics, the tunneling effect describes the penetration of particles through barriers, even if the particle’s energy is less than the barrier height., which allows it to potentially surpass this higher barrier directly and thus reach the minimum faster [6]. Recent advancements in QA methods have suggested potential benefits by employing QA over SA to address NP-hard optimization challenges [18].
@@ 2.4 Binary Quadratic Models
For QA, cost functions are mathematical expressions that are usually formulated as binary quadratic models666https://docs.dwavequantum.com/en/latest/quantum_research/qubo_ising.html and whose lowest values generally represent the best solutions to problems [36, 38]. To solve problems on the D-Wave quantum annealer, the problem space has to be formulated as a BQM, the D-Wave Systems software representation of the objective function [38]:
where s and x are binary variables, i.e., categorical variables that can only take one of two values. s is the decision variable called spin that can take values from the set {-1,1}, in which case the problem is called an Ising problem (Eq. 2) [15]. In the Ising model, hi is the magnetic field that represents the linear coefficients corresponding to qubit biases and Ji,j represents the coupling or interaction strength between qubits [38, 36]. In a QUBO problem (Eq. 3), x can take values from the set {0,1} [13]. Qubit biases are represented by the linear coefficients ai and the coupling strength between qubits by the quadratic coefficients bi,j [30]. Both versions are NP-hard and can easily be converted into each other by applying variable substitution [30, 34]. In this paper, the QUBO format is used to formulate the problem. This is because QUBO formulations are a widely used method for this [9].
The QUBO formulation is further mapped to a QPU topology [36] to solve the problem during the annealing process and minimize the energy function. QA devices are less affected by noise than gate model quantum computers. This allows for more qubit usage and thus solves larger problems. QA provides an approach that promises good solutions to NP-hard problems and specializes in combinatorial optimization. While the discussion continues whether QA can provide quantum speed-up over classical approaches, there have been several attempts to solve application-motivated problems with QA [36].
@@ 3 Related Work
In this Section, we assess previous approaches to solving the optimization problems described in Subsection 2.1.
@@
JSSP. In the past, a variety of approaches to solve the JSSP or its variation existed. Denkena et al. [7] mention heuristics, Genetic Algorithm (GA), Tabu Search, Ant Colony, Particle Swarm Optimization and SA. These approaches are used both individually and in hybrid combinations with one another. Venturelli et al. [33] present a solver for the JSSP that makes use of a QA. Carugno et al. [5] studied the application of QA to solve JSSP and discussed challenges related to problem formulation for the fine-tuning of the QA. In terms of solution quality, QA is remarkably effective in comparison to classical approaches on very constrained problem instances, while closely competitive on realistic problem instances. Pakhomchik et al. [22] focus on the mathematical formulation of JSSP and propose a new large neighborhood search approach, with good empirical results for large problem instances.
Zielewski et al. [40] use QA to solve JSSP and especially investigate the embedding process in the performance. The authors demonstrate that the initial setup of the problem, post-processing, and problem formulation, as well as the presence of excess variables and large embeddings, have a significant impact on the results. A. Sharma and P. Maharjan [2] compare QA to SA through existing published works and conclude that QA has the potential to be considered as an optimized technique against SA. QA is more efficient than SA in estimating the ground state of the system for simulations of small and large sample sizes. Denchev et al. [6] show that, for problems where local minima is separated by tall and narrow energy barriers, QA is more than times faster than SA running on a single core. Wei et al. [35] introduce a hybrid quantum-classical Benders’ decomposition for federated learning scheduling, reducing iteration counts by 70.3% and computation time by 81%. Smierzchalski et al. [28] show that a hybrid quantum-classical computation scheme can be efficiently used for automatic guided vehicle scheduling in manufacturing.
@@
In general, research considering transportation time focuses on the flow shop scheduling problem777Flow shop scheduling is a special case of job shop scheduling in which all jobs have identical operation flow patterns on machines [1]. or classical FJSP. Most of the QA research in this field focuses on solving JSSP or FJSP, and transportation time is often neglected in the literature. However, transportation time between locations can have a significant impact on the maximum completion time of FJSP.
@@
FSP. Flow Shop Scheduling Problem (FSP) is one of the most studied problems in order planning. Hurink and Knust [10] apply SA to minimize both the total completion time and the total tardiness of flow shop problems. Naderi et al. [19] propose an adaptation of the SA algorithm to solve the flow shop problem with incorporated sequence-dependent transportation times with a single-transporter. Several heuristic algorithms were developed considering the transportation of an interrupted job from one machine to another. Naderi et al. [20] consider the transportation times in permutation flow shops and provide solution methods that include heuristics and metaheuristics to solve large problems. An adaptation of the imperialist competitive algorithm, hybridized with a simulated annealing-based local search, was proposed by Karimi et al. [12] to solve the problem.
@@
FJSP. Schworm et al. [25] investigate a QA-based FJSP on large problem instances specific to modern manufacturing and demonstrate the efficiency of the approach regarding scalability, solutions’ quality, and computing time. Denkena et al. [7] compare the computing time of several approaches to solve the classical FJSP and conclude that it generally increases significantly with the size of the problem. Their research suggests that these methods deliver valid results for small problem instances, but are not suitable for problem sizes relevant to practical applications. Therefore, methods are required that yield high-quality outcomes while also ensuring consistently quick computation times, particularly for large problem instances. The paper also investigates the concept of QA- based optimization for a process-parallel FJSP on existing benchmarks. The results reveal that the QA technique consistently achieves low anneal times, especially for large instances, while still producing outcomes comparable to those of classical heuristics. Recent work by Fu et al. [8] solves the FJSP using a Coherent Ising Machine (CIM) and shows an 80% improvement compared to classical methods, but only for small instances, due to existing hardware constraints. INOUE et al. [11] analyzed the FJSP with Tool Switching Constraints with the Constrained Quadratic Model (CQM) solver supported by the D-Wave Leap Hybrid Solver and showed that the quality and computation time significantly outperform the results achieved using Dispatching rules or Python Mixed-Integer Linear programs. Schworm et al. [26] addressed real-time responsiveness to disturbances in manufacturing environments using QA for dynamic job shop scheduling and showed an advantage for large problem sizes. Amaro et al. [3] present a case study applying variational quantum algorithms to job shop scheduling with notably accelerated convergence. Recent results by Tang et al. [31] show that quantum computing can deliver dramatic efficiency gains for several automated guided vehicles scheduling models in manufacturing, with up to 92% time savings. Ren and Liu [24] introduce a dynamic FJSP approach that incorporates machine breakdowns and new job arrivals, thereby furthering the practicality of quantum-inspired scheduling.
DFJSP. Only few papers focus on the FJSP with transportation time between jobs, also known as DFJSP. Marzouki et al. [16] developed a novel mathematical model for the distributed DFJSP that explicitly incorporates transportation time between machines, and proposed an effective optimization framework that significantly reduces makespan and enhances coordination across distributed manufacturing environments. Peng et al. [23] formulated a multi-objective dynamic DFJSP that explicitly accounts for uncertain processing times, proposing a novel optimization framework that improves schedule robustness and adaptability in distributed manufacturing environments under real-world uncertainty. Yuan et al. [37] proposed an improved deep Q network approach for the distributed heterogeneous FJSP that integrates automated guided vehicle (AGV) transportation, demonstrating substantial optimization of production makespan and resource utilization. Zhang et al. [39] develop a modified Genetic Algorithm (GA) to solve the FJSP with transportation time. Marzouki et al. [17] use a Chemical Reaction Optimization Metaheuristic to solve the DFJSP. Liu et al. [14] propose a refined GA and apply it with satisfactory results on a real case of a manufacturing company.
@@
Generally, most of the papers focus on studying the JSSP. Although some research includes transportation time in the model, these studies typically employ classical approaches, such as heuristics or GA. In this research, an extended case of DFJSP is studied, where not only jobs (production orders), but also job operations (production steps) can be distributed. For this purpose, a D-Wave Quantum Annealer is used to solve the DFJSP with distributed operations. The availability of a set of machines capable of processing the same operation and the geographical distribution of these machines add complexity and pose a significant challenge to solve this combinatorial optimization problem.